# Random large eddy simulation for 3-dimensional incompressible viscous flows

### v1

@@ 1 Introduction
The idea of utilising statistical mechanics and probabilistic approach is not new in Computational Fluid Dynamics (CFD), although, to the best knowledge of the present authors, it seems that methods such as Monte-Carlo simulation, particle system method of turbulence and other stochastic methods have not been used extensively in industry. While these methods are quite stimulating and there is a high potential of being useful tools in the study of fluid flows including turbulent flow simulation. Among them, the random vortex method (cf. [1], [2] and [8] for details) developed over the last three decades stands out as an important tool, whose successful applications in simulations of incompressible flows however are limited to special flows, mainly are restricted to two dimensional flows. The random vortex method for three dimensional (3D) flows has been developed recently in [11] by using a Feynman-Kac type formula, revealing the difficulty caused by the non-linear stretching term which does not appear in two dimensional flows. Technically, the main problem in the random vortex method for three dimensional incompressible flows lies in the fact the singular integral kernel appearing in the stochastic integral representation is no longer locally integrable, which certainly induces instability of corresponding numerical schemes. Let us recall the key idea in random vortex methods. Let $u(x,t)$ and $p(x,t)$ denote the velocity and the pressure respectively. Then the motion equation of the flow is the Navier-Stokes equations (cf. [5]):
@@
where $\nu>0$ is the kinematic viscosity, and $F$ is an external force applying to the flow. The key observation is that the velocity $u$ can be recovered from the vorticity $\omega=\nabla\wedge u$ through the Biot-Savart law by integrating the vorticity against a singular integral kernel $K(y,x)$. In fact, $K$ is the gradient of the Green function, and therefore $K$ is a nice kernel for integration. While for three dimensional flows, a further derivative of the Biot-Savart kernel $K$ is required, which is not locally integrable unfortunately. This difficulty originates from the random vortex method which aims to calculate numerically the vorticity $\omega$ first based on the vorticity transport equation
@@
and the velocity $u$ is recovered from $\omega$. Let us recall the technical steps briefly for two dimensional flows and the external vorticity $\nabla\wedge F=0$ to motivate our approach in the present paper. For this case, the non-linear stretching term $(\omega\cdot\nabla)u$ vanishes identically. Let $X$ be Brownian fluid particles $X$ with velocity $u(x,t)$. That is, according to Taylor’s Brownian motion $X$ is determined by stochastic differential equation
@@
To avoiding confusion, the notation $X^{\xi}$ denotes Brownian fluid particles issued from a location $\xi$. Since $\omega$ is a solution of the vorticity transport equation, an easy exercise shows that
@@
$\textrm{d}X_{t}^{\xi}=\left(\int\left.\mathbb{E}\left[K(X_{t}^{\eta},x)\right]u_{0}(\eta)\textrm{d}\eta\right|_{x=X_{t}^{\xi}}\right)\textrm{d}t+\sqrt{2\nu}\textrm{d}B,\quad X_{0}^{\xi}=\xi$
on which numerical schemes may be developed accordingly for calculating two dimensional flows. The idea similarly applies to three dimensional flows but unfortunately involves the gradient of $K$ which is not locally integrable. See more details about random vortex method and integral representation theorem in [6], [11] and [12].
There are mainly two difficulties in numerically solving fluid dynamic equations of turbulent flows, which are non-linear characteristics of the Navier-Stokes equation and continuously changing complex vortices of different scales. In engineering applications, people are more concerned about the time-averaged effect of turbulent motion, so some of the commonly used turbulence models are based on Reynolds time averaging, called Reynolds Average Numerical Simulation (RANS). RANS smooths out some tiny details of turbulent motion, and the model has many artificial settings. Therefore, RANS models have limited simulation capabilities for complex and delicate turbulent structure such as the separation phenomena of flow past body. With the increasing of the computing power and computer memory capacity, some advanced research institutions solve the Navier-Stokes equations without any form of simplification, but use extremely fine grids to directly numerically solve the Navier-Stokes equations. This kind of approach may be called Direct Numerical Simulation (DNS). However, DNS are still computational expanse to implement. The Large Eddy Simulation (LES) method, which is between DNS and the RANS method, has gradually emerged in the computational fluid dynamics community and has developed into a promising method for numerically solving turbulent flows because it is more sophisticated than the RANS and can be implemented on a conventional computer.
@@
where $x,x^{\prime}\in\mathbb{R}^{3}$, $\chi(x,t)$ is the filtering function including top-hat function, Gaussian function and so on, which satisfies $\intop\chi(x)\mathrm{d}x=1$. So the filtered Navier-Stokes equation becomes
@@
Because $\widetilde{(u\cdot\nabla)u}$ can not be simplified as $(\tilde{u}\cdot\nabla)\tilde{u}$, so researcher introduced sub-grid-scale stress $\tau$, defined as
@@
The $\tau$ means the effect of small eddies on the flow system which must be modelled because its information is not irrelevant. The influence of small-scale eddies on the equation of motion is described by some other models (cf. [9], [14] and [10]), but it is not the studying target of our paper. Through large eddy simulation, the filtered equations focus on analysing large-scale eddies, directly capturing the most important turbulent structure and dynamical information, while approximating small-scale eddies. This method effectively reduces the computational complexity and retains the main physical properties.
Random large eddy simulation method we developed in this paper combines the core idea of filtering in large eddy simulation with the integral representation of parabolic equation in random vortex method, and subsequently obtain the numerical method of incompressible viscous flows based on it. Using the stochastic integral representation of parabolic equation and basic idea of large eddy simulation, we are able to close the solution of Navier-Stokes equation, which makes it possible for us to simulate incompressible flows via Monte-Carlo method.
The paper is organised as follows. In Section 2, we propose the random large eddy simulation method based on stochastic integral representation and large eddy simulation. Then, numerical schemes and simulation experiments are presented in Section 3. Finally, we introduce conditional law duality and integral representation theorem in Appendix, which is the theoretical foundation of the random large eddy simulation method established in Section 2.
@@ 2 Random large eddy simulation method
by using the formal symbol $h(0,\eta;t,x)=\mathbb{P}\left[X_{t}^{\eta}=x\right]$. The problem here is that of course $\delta_{x}(X_{t}^{\eta})$ is a generalised Wiener functional so that it is difficult to calculate it numerically. Also the conditional average $\mathbb{E}\left[\left.\delta_{x}(X_{t}^{\eta})\right|X_{s}^{\eta}\right]$ is expensive to compute. To overcome these difficulties, we utilise a key idea from the approach of Large Eddy Simulations (LES). In LES, one calculates local averaged velocity which shall give the global structure of the flow (in particular for turbulent flows). We shall adopt the LES approach, avoiding the modelling of the error term caused by taking local average of the velocity.

### v2

@@ 1 Introduction
The idea of utilising probabilistic approach is not new in Computational Fluid Dynamics (CFD), although, to the best knowledge of the present authors, it seems that methods such as Monte-Carlo simulation, particle system method of turbulence and other stochastic methods have not been used extensively in industry. While these methods are quite stimulating and there is a high potential of being useful tools in the study of fluid flows including turbulent flow simulation. Among them, the random vortex method (cf. [1], [2] and [8] for details) developed over the last three decades stands out as an important tool, whose successful applications in simulations of incompressible flows however are limited to special flows, mainly are restricted to two dimensional flows. The random vortex method for three dimensional (3D) flows has been developed recently in [11] by using a Feynman-Kac type formula, revealing the difficulty caused by the non-linear stretching term which does not appear in two dimensional flows. Technically, the main problem in the random vortex method for three dimensional incompressible flows lies in the fact the singular integral kernel appearing in the stochastic integral representation is no longer locally integrable, which certainly induces instability of corresponding numerical schemes. Let us recall the key idea in random vortex methods. Let $u(x,t)$ and $p(x,t)$ denote the velocity and the pressure respectively. Then the motion equation of the flow is the Navier-Stokes equations (cf. [5]):
@@
where $\nu>0$ is the kinematic viscosity, and $F$ is an external force applying to the flow. The key observation is that the velocity $u$ can be recovered from the vorticity $\omega=\nabla\wedge u$ through the Biot-Savart law by integrating the vorticity against a singular integral kernel $K(y,x)$. In fact, $K$ is the gradient of the Green function, and therefore $K$ is a nice kernel for integration. While for three dimensional flows, a further derivative of the Biot-Savart kernel $K$ is required, which is not locally integrable unfortunately. This difficulty originates from the random vortex method which aims to calculating numerically the vorticity $\omega$ first based on the vorticity transport equation
@@
and the velocity $u$ is recovered from $\omega$. Let us recall the technical steps briefly for two dimensional flows and the external vorticity $\nabla\wedge F=0$ for motivating our approach in the present paper. For this case, the non-linear stretching term $(\omega\cdot\nabla)u$ vanishes identically. Let $X$ be Brownian fluid particles $X$ with velocity $u(x,t)$. That is, $X$ is determined by stochastic differential equation
@@
For avoiding confusion, we use $X^{\xi}$ to denote Brownian fluid particles issued from a location $\xi$. Since $\omega$ is a solution of the vorticity transport equation, an easy exercise shows that
@@
$\textrm{d}X_{t}^{\xi}=\left(\int\left.\mathbb{E}\left[K(X_{t}^{\eta},x)\right]u_{0}(\eta)\textrm{d}\eta\right|_{x=X_{t}^{\xi}}\right)\textrm{d}t+\sqrt{2\nu}\textrm{d}B,\quad X_{0}^{\xi}=\xi,$
and therefore numerical schemes may be developed accordingly for calculating two dimensional flows. The idea similarly applies to three dimensional flows but unfortunately the resulting stochastic integral equation involves the gradient of $K$, cf. [6], [11] and [12] for the details about random vortex method and integral representation theorem.
There are mainly two difficulties in numerically solving fluid dynamic equations of turbulent flows, which are non-linear characteristics of the Navier-Stokes equation and continuously changing complex vortices of different scales. In engineering applications, people are more concerned about the time-averaged effect of turbulent motion, so some of the commonly used turbulence models are based on Reynolds time averaging, called Reynolds Average Numerical Simulation (RANS). RANS smooths out some tiny details of turbulent motion, and the model has many artificial settings. Therefore, RANS models have limited simulation capabilities for complex and delicate turbulent structure such as the separation phenomena of flow past body. With the increasing of the computing power and computer memory capacity, some advanced research institutions solve the Navier-Stokes equations without any form of simplification, but use extremely fine grids for directly numerically computing solutions of the Navier-Stokes equations. This kind of approach may be called Direct Numerical Simulation (DNS). However, DNS are still computational expanse to implement. The Large Eddy Simulation (LES) method, which is between DNS and the RANS method, has gradually emerged in CFD community and has developed into a promising method for numerically solving turbulent flows because it is more sophisticated than the RANS and can be implemented on a conventional computer.
@@
where $x,x^{\prime}\in\mathbb{R}^{3}$, $\chi(x,t)$ is the filtering function including top-hat function, Gaussian function and so on, which satisfies that $\intop\chi(x)\mathrm{d}x=1$. So the filtered Navier-Stokes equation becomes
@@
Because $\widetilde{(u\cdot\nabla)u}$ can not be simplified as $(\tilde{u}\cdot\nabla)\tilde{u}$, so researchers introduced sub-grid-scale stress $\tau$, defined as
@@
The tensor $\tau$ measures the effect of small eddies on the flow system which must be modelled because its information is not irrelevant. The influence of small-scale eddies on the equation of motion is described by some other models (cf. [9], [14] and [10]), but it is not the study target of our paper. Through LES, the filtered equations focus on analysing large scale eddies, directly capturing the most important turbulent structure and dynamical information, while approximating small-scale eddies. This method effectively reduces the computational complexity and retains the main physical properties.
Random large eddy simulation method we are going to develop in the present paper combines the core idea of filtering in LES with the integral representation of parabolic equation in random vortex method, and subsequently obtain the numerical method of incompressible viscous flows based on it. Using the stochastic integral representation of parabolic equation and basic idea of LES, we are able to close the solution of Navier-Stokes equation, which makes it possible for us to simulate incompressible flows via Monte-Carlo method.
The paper is organised as follows. In Section 2, we propose the random large eddy simulation method based on stochastic integral representation and LES. Then, numerical schemes and simulation experiments are presented in Section 3. Finally, we introduce conditional law duality and integral representation theorem in Appendix, which is the theoretical foundation of the random large eddy simulation method established in Section 2.
@@ 2 Random large eddy simulation method
by using the formal symbol $h(0,\eta;t,x)=\mathbb{P}\left[X_{t}^{\eta}=x\right]$. The problem here is that of course $\delta_{x}(X_{t}^{\eta})$ is a generalised Wiener functional so that it is difficult to calculate it numerically. Also the conditional average $\mathbb{E}\left[\left.\delta_{x}(X_{t}^{\eta})\right|X_{s}^{\eta}\right]$ is expensive to compute. To overcome these difficulties, we utilise a key idea from LES. In LES, one calculates local averaged velocity which shall give the global structure of the flow (in particular for turbulent flows). We shall adopt the LES approach, avoiding the modelling of the error term caused by taking local average of the velocity.